OpenStudy (anonymous):

What is a solution for the system of equations y=2x-15 and y=x^2-6x? thank you.

OpenStudy (wach):

y=2x-15, y=x^2-6x? I assume you're trying to find a value that works for both equations? If so, you first want to make them equal, as y variable is able to link them: So, 2x-15=x^2-6x subtract -2x from both sides, add 15 to both sides 0=x^2-8x+15 Solve using squares (if you are familiar with this) 0= (x-5)(x-3) Thus, your answer set for both is (5) and (3) If that helps :)

OpenStudy (anonymous):

that was very helpful thank you!